A 1.50 L buffer solution is 0.250 M in HF and0.250
Answer #1

Ka of HF = 3.5*10^-4
pka = -log Ka
= -log (1.52*10^-5)
= 3.46

mol of HF present initially = M*V = 0.250 M * 1.5 L = 0.375 mol
mol of NaF present initially = 0.250* 1.5 =0.375 mol

mol of NaOH added = 0.05 mol

0.05 mol of HF and NaOH will react to form extra 0.05 mol of NaF

after reaction,
mol of NaF= 0.375 + 0.05 = 0.425 mol
mol of HF remaning = 0.375 - 0.05 mol = 0.325 mol

total volume = 1.5 L
[NaF] = 0.425/1.5 = 0.283 M
[HF] = 0.325/1.5 = 0.217 M

use:
pH = pKa + log {[NaF]/[HF]}
= 3.46 + log (0.283/0.217)
= 3.57
Answer: B